$SPAD/src/input f07fdf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
uplo:="L"
n:=4
lda:=4
a:Matrix SF:=
 [[4.16 ,0.0 ,0.0 ,0.0 ],_
  [-3.12 ,5.03 ,0.0 ,0.0 ],_
  [0.56 ,-0.83 ,0.76 ,0.0 ],_
  [-0.1 ,1.18 ,0.34 ,1.18 ]]
result:=f07fdf(uplo,n,lda,a)

———-

References^{[1] nothing}

4